OpenStudy (anonymous):

Use the Single Tax Brackets chart to answer the question that follows. Single Payer Tax Brackets Taxable Income: Income Tax: $0–$8,375 10% of the amount $8,375–$34,000 $837.50 plus 15% of the amount over $8,375 $34,000–$82,400 $4,681.25 plus 25% of the amount over $34,000 $82,400–$171,850 $16,781.25 plus 28% of the amount over $82,400 $171,850–$373,650 $41,827.25 plus 33% of the amount over $171,850 $373,650+ $108,421.25 plus 35% of the amount over $373,650 James has a taxable income of $167,391. What is his tax liability?

OpenStudy (anonymous):

@jim_thompson5910 this one?

jimthompson5910 (jim_thompson5910):

unfortunately taxes aren't as simple as taking a single percentage and you're done that would be nice though

OpenStudy (anonymous):

fyi, there are forty questions, i only need help on like 5 that i completely dont know

jimthompson5910 (jim_thompson5910):

what you do is start with the lowest amount, evaluate the tax on that then move up the bracket and you add on taxes based on the amount over the tax bracket thresholds

OpenStudy (anonymous):

see, this is why im asking my peers, because i dont know any of the questions im asking

jimthompson5910 (jim_thompson5910):

the taxable income is 167,391 the first 8,375 is taxed at 10% so you take 10% of 8,375 to get 837.50 ---------------------------------------- now you move onto the next tax bracket and you only apply the tax rate here for the remaining amount over 8,375 so you are taxing the remaining 167,391 - 8,375 = 159,016 at 15% and adding it to the total tax amount see how this is all working out?

jimthompson5910 (jim_thompson5910):

luckily, they seemed to have done much of the work for us because I just noticed it says "$837.50 plus 15% of the amount over $8,375" so they've already computed the lower tax bracket amounts

OpenStudy (anonymous):

No sir, im sorry

jimthompson5910 (jim_thompson5910):

hmm well taxes are definitely confusing (much of it is on purpose sadly)

OpenStudy (anonymous):

I can tell, my brain is not liking taxes

jimthompson5910 (jim_thompson5910):

the important thing to notice is that $167,391 is in the $82,400–$171,850 tax bracket

jimthompson5910 (jim_thompson5910):

so you apply the rule $16,781.25 plus 28% of the amount over $82,400

jimthompson5910 (jim_thompson5910):

so no matter what, you're paying $16,781.25

OpenStudy (anonymous):

I understand that because its common sense haha

jimthompson5910 (jim_thompson5910):

you then add on 28% of what's left over (after subtracting off 82,400) so 16,781.25 - 82,400 = 84,991 then you apply the 28% rate to this portion to get 0.28*84,991 = 23,797.48 so you add on this amount to the 16 grand amount we already stated above 16,781.25 + 23,797.48 = 40,578.73

OpenStudy (anonymous):

You are typing something biiiiiiiiiiiiiiiiiiiiiiiiiiig

jimthompson5910 (jim_thompson5910):

or just something with a lot of calculations

OpenStudy (anonymous):

true, dang son, you are a mathematician, its crazy

jimthompson5910 (jim_thompson5910):

so he owes $40,578.73 in taxes

jimthompson5910 (jim_thompson5910):

nah this is just a lot of arithmetic (and crazy hard to remember rules)

OpenStudy (anonymous):

Well, thanks again buddy, im about to ask another few questions haha, and i guess, i thought i used to be good at math... jokes

jimthompson5910 (jim_thompson5910):

you just need more practice and if you've never seen stuff like this before, then it's going to be crazy

OpenStudy (anonymous):

for sure